# Lyapunov exponents and geodesic stability of Schwarzschild black hole in the non-commutative gauge theory of gravity

### v3

@@ III.4.1 Null circular orbit and photon sphere
The conditions for the circular motion of photons in NC Schwarzschild spacetime are analogous to those for massive particles, as given in (33). By utilizing the effective potential (23) with $h=0$, the relationship between the photon’s energy $E_{c}$ and angular momentum $L_{c}$ can be expressed as:
$\frac{1}{D_{c}}=\frac{E_{c}}{L_{c}}=\sqrt{\frac{r_{c}-2m}{r_{c}^{3}}+\left(\frac{J(r_{c})+G(r_{c})\sqrt{1-\frac{2m}{r_{c}}}}{16r_{c}^{6}(r_{c}-2m)}\right)\Theta^{2}},$ (51)
@@
$J(r_{c})=156m^{3}-185m^{2}r_{c}+73mr_{c}^{2}-10r_{c}^{3},\quad G(r_{c})=33m^{2}r_{c}-37mr_{c}^{2}+6r_{c}^{3}.$ (52)
Table 1: Numerical values for the unstable $r_{\text{uns}}$ and the new stable circular orbit $r_{\text{sta}}$ of the photon sphere in NC spacetime for different $\Theta$ values, with $E=0.998$, $l=4.2$, and $m=1$.
@@
$r_{\text{sta}}$ (new) 2.18400 2.24624 2.30383 2.35865 2.41183
$r_{\text{uns}}$ 3.00000 2.99782 2.99504 2.99101 2.98559 2.97855
Table 1 presents numerical solutions of (33), illustrating the dependence of the stable and unstable photon sphere radii on $\Theta$. For $3.29\leq l\leq 4.48$ (where this condition depends on $0\leq\Theta\leq 0.3$), the stable circular orbit radius increases with $\Theta$, while the unstable orbit radius decreases. However, for $l>4.48$, both orbit radii increase as $\Theta$ grows.
Figure 11: Behavior of the photon sphere radius in NC spacetime. Stable and unstable circular orbits are shown as functions of $\Theta$ for fixed $l=4.2$, $E=1$, $m=1$ (two left panels), and as functions of $l$ for fixed $\Theta=0.3$, $E=1$, $m=1$ (two right panels).
The left panels of Fig. 11 demonstrate that as $\Theta$ increases, the radius of the stable circular orbit grows, while the radius of the unstable orbit decreases. In the NC framework, the unstable photon orbit has a smaller radius than in the commutative case for $l<4.48$, whereas for $l\geq 4.48$, the unstable orbit radius exceeds the commutative value. This divergence arises from modifications to the gravitational field induced by non-commutativity. The right panels of Fig. 11 illustrate the dependence of circular orbit radii on angular momentum $l$. For a fixed $\Theta=0.3$, the stable circular orbit radius decreases as $l$ increases, while the unstable orbit radius grows. This behavior reflects the distinctive interplay between angular momentum and the NC parameter in shaping the geodesic structure of spacetime.
@@ III.4.2 Lyapunov exponents
In the following, we analyze the stability and instability of the two photon spheres in NC spacetime using the Lyapunov exponent. As is well known, for null circular orbits of photons, proper time does not exist, leaving only the coordinate time Lyapunov exponent for analysis.
We begin by calculating the angular frequency and the period in coordinate time, as given by equations (37) and (40), respectively:
@@
Next, the coordinate time Lyapunov exponent for photon orbits around the NC Schwarzschild BH is derived as:
$\hat{\lambda}_{c}^{\text{Null}}=\sqrt{-\frac{3((-4m+r_{c})(-2m+r_{c}))}{r_{c}^{4}}+\left(\frac{O(r_{c})+I(r_{c})\sqrt{1-\frac{2m}{r_{c}}}}{32(r_{c}-2m)^{2}r_{c}^{7}}\right)\Theta^{2}},$ (54)
@@
Fig. 13 illustrates the behavior of the coordinate time Lyapunov exponent for photons as a function of $r_{c}$ (left panel) and $\Theta$ (right panel). The left panel shows that instability initially increases to a peak (decreasing with higher $\Theta$), then decreases as $r_{c}$ increases, eventually vanishing for $r_{c}\geq 4$. In NC spacetime, orbits near the event horizon exhibit reduced instability compared to the commutative case, whereas at larger distances from the event horizon, this trend reverses. Moreover, while instability begins at the event horizon in commutative spacetime, NC geometry shifts it outward, allowing for the presence of stable orbits near the event horizon. This observation is consistent with the effective potential analysis for massless test particles, confirming the existence of a stable photon sphere around the NC Schwarzschild BH. Similar behavior has been reported in other NC geometries, as discussed in Ref. geostab5 , where orbits within the event horizon cannot form a stable photon sphere. Examining the dependence of circular orbit radii $r_{c}$ on $\Theta$ (right panel) reveals that the orbit at $r_{c}=3$ remains unstable in NC spacetime, with increasing $\Theta$ amplifying this instability. In contrast, the new circular orbits emerging near the event horizon in NC spacetime exhibit reduced instability, which decreases as $\Theta$ increases and eventually stabilizes. This suggests the existence of stable circular orbits near the event horizon in NC spacetime.
The ratio of the Lyapunov exponent to the angular frequency in coordinate time ($\frac{\hat{\lambda}_{c}^{\text{Null}}}{\hat{\Omega}_{c}^{\text{Null}}}$) defines the instability exponent of null circular orbits in NC spacetime:
$\frac{\hat{\lambda}_{c}^{\text{Null}}}{\hat{\Omega}_{c}^{\text{Null}}}=\sqrt{\frac{3(4m-r_{c})}{r_{c}}+\left(\frac{O(r_{c})+6J(r_{c})(8m^{2}-6mr_{c}+r_{c}^{2})+(I(r_{c})+6G(r_{c})(8m^{2}-6mr_{c}+r_{c}^{2}))\sqrt{1-\frac{2m}{r_{c}}}}{32(r_{c}-2m)^{3}r_{c}^{4}}\right)\Theta^{2}}.$ (56)
@@
Fig. 14 illustrates the variation of $\hat{\lambda}_{c}^{\text{Null}}/\hat{\Omega}_{c}^{\text{Null}}$ with $r_{c}$, showing behavior similar to that in Fig. 13. Unlike the commutative case, NC spacetime allows for stable orbits near the event horizon, in contrast to the results in Ref. geostab5 , which confirm the stability of the new photon sphere.
Finally, the critical exponent $\hat{\gamma}$ for null circular orbits is defined as the ratio of the Lyapunov timescale $\hat{T}_{\hat{\lambda}}=\frac{2\pi}{\hat{\lambda}}$ to the orbital timescale $\hat{T}_{\hat{\Omega}}=\frac{2\pi}{\hat{\Omega}}$ geostab11 ; geostab3 ; geostab12 :
$\hat{\gamma}^{\text{Null}}=\frac{\hat{T}_{\hat{\lambda}}^{\text{Null}}}{\hat{T}_{\hat{\Omega}}^{\text{Null}}}=\frac{1}{2\pi\sqrt{\frac{3(4m-r_{c})}{r_{c}}+\left(\frac{O(r_{c})+6J(r_{c})(8m^{2}-6mr_{c}+r_{c}^{2})+(I(r_{c})+6G(r_{c})(8m^{2}-6mr_{c}+r_{c}^{2}))\sqrt{1-\frac{2m}{r_{c}}}}{32(r_{c}-2m)^{3}r_{c}^{4}}\right)\Theta^{2}}}.$ (57)
@@
The numerical values of the critical exponent $\hat{\gamma}^{\text{Null}}$ for photons in various circular orbits $r_{c}$ and different values of the non-commutative (NC) parameter $\Theta$ are summarized in Table 2. This table highlights the presence of both real and imaginary values of $\hat{\gamma}^{\text{Null}}$. Real positive values indicate unstable circular orbits geostab11 , whereas imaginary values correspond to stable circular orbits. For the photon sphere orbit in the Schwarzschild BH spacetime at $r_{c}=3$, instability persists even in the NC framework. According to the definition in Eq. (57), this implies that the Lyapunov timescale $\hat{T}_{\hat{\lambda}}$ is shorter than the orbital timescale $\hat{T}_{\hat{\Omega}}$, i.e., $\hat{T}_{\hat{\lambda}}<\hat{T}_{\hat{\Omega}}$ geostab12 . A similar behavior is observed for other values of $r_{c}$ in Table 2, where real values of $\hat{\gamma}^{\text{Null}}$ consistently signify instability. Furthermore, the impact of the NC parameter $\Theta$ on $\hat{\gamma}^{\text{Null}}$ is evident. As $\Theta$ increases, the values of $\hat{\gamma}^{\text{Null}}$ decrease for $r_{c}=3$, indicating increasing instability. The smallest values of $\hat{\gamma}^{\text{Null}}$ correspond to stronger Lyapunov instability geostab3 (see the right panel of Fig. 13). On the other hand, the imaginary values of $\hat{\gamma}^{\text{Null}}$, which denote stable circular orbits, align with the new minima in the effective potential observed for $l=5.2$ (values comparable to those in Table 1). This demonstrates the emergence of a new stable photon sphere in NC spacetime, contrasting with the instability of the external photon sphere in both commutative and NC scenarios.
III.4.3 Black hole shadow
@@
In the equatorial plane ($\theta=\pi/2$), the radius of the BH shadow, $\hat{R}_{\text{shadow}}$, is directly related to the impact parameter $D_{c}$ by the relation $\hat{R}_{\text{shadow}}=D_{c}\big{|}_{r=r_{\text{ps}}}$, where $r^{\text{uns}}_{\text{ps}}$ represents the radius of the unstable photon sphere. This radius can be determined by solving the equation $\frac{dV_{\text{eff}}}{dr}=0$. To leading order in the NC parameter $\Theta$ and the BH mass $m$, the solution is given by:
$r^{\text{uns}}_{\text{ps}}=3m-\left(\frac{-38+540D_{c}^{2}m^{2}}{288m}\right)\Theta^{2}\,.$ (58)
@@
It is important to note that the radius of the stable photon sphere in NC spacetime can only be obtained numerically (see Table 2). Using the above expression, the shadow radius of the NC Schwarzschild BH can be expressed as:
@@
$\hat{R}_{\text{shadow}}=3\sqrt{3}m\left(1+\frac{(1+2\sqrt{3})}{72m^{2}}\Theta^{2}\right)\,,$ (59)
@@
Fig. 15 illustrates the shadow of the NC Schwarzschild BH for different values of the NC parameter $\Theta$ (left panel) and the BH mass $m$ (right panel). As seen in the left panel, the shadow radius increases with increasing $\Theta$. Similarly, the right panel shows that for a fixed NC parameter, the shadow radius grows as the BH mass $m$ increases. This behavior suggests that the effects of non-commutativity and BH mass on the shadow radius are analogous. The NC parameter influences the spacetime geometry in a way that mimics an increase in mass, effectively strengthening the gravitational field. This conclusion is consistent with our earlier results abdellah1 . Furthermore, we observe that the shadow radius of the NC Schwarzschild BH is larger and exhibits a slight increase with increasing $\Theta$, in contrast to the findings in Hassanabadi2 , where it was reported to increase significantly with increasing $\Theta$.
@@
Armed with the above result and considering the constraint on the NC parameter—namely, that the NC correction must be smaller than the precision of the experimental measurement we choose as an implementation of the NC shadow radius of the Schwarzschild BH. Specifically, we consider a typical primordial BH at the final stage of inflation with a mass of $GM\sim 5\times 10^{-5}m$ karimabadi1 and a volume of $r\sim 1.5\times 10^{-4}m$. Moreover, the precision of the experimental measurement, according to ehc , is approximately $\sim 0.1$. Consequently, the upper limit of the NC parameter can be determined as:
$\Bigg{|}\frac{R_{\text{shadow}}-\hat{R}_{\text{shadow}}}{R_{\text{shadow}}}\Bigg{|}=\sqrt{3}\frac{(1+2\sqrt{3})}{24m}\Theta^{2}\leq 0.1\,.$ (60)
@@
According to Ref. PLANCK1 , the physical distance measured at any time after inflation can be obtained by multiplying our results by the square of the cosmological scale factor, $a=10^{-29}$. This is due to our choice of the space-space NC matrix. By applying the above constraint to our system, we find that the NC parameter satisfies:
$\Theta^{\text{Phy}}=\sqrt{a^{2}\Theta^{2}}\leq 3.94\times 10^{-32}\,\text{m}.$ (61)
@@
It is evident that the obtained result is close to the Planck scale and is consistent with the bounds derived in our previous works, where similar constraints were obtained through experimental tests of gravity in the context of the NC gauge theory of gravity abdellah1 ; abdellah4 or by employing the Bopp’s shift transformation larbi1 ; mirza ; rome1 . These studies suggest a bound in the range of $(10^{-34}-10^{-31})$ m. Remarkably, in this work, the estimation of the NC parameter is directly related to the accuracy of the experimental measurement of the BH shadow radius. This implies that as the precision of the measurement improves, the estimation of the NC parameter can become more refined. A similar observation was made in Ref. abdellah4 , where variations in experimental accuracy led to different values of the NC parameter.

### v4

@@ III.4.1 Null circular orbit and photon sphere
The conditions for circular photon orbits in NC Schwarzschild spacetime are analogous to those for massive particles, as given in Eq. (32). By utilizing the effective potential (22) with $h=0$, the relationship between the photon’s energy $E_{c}$ and angular momentum $L_{c}$ can be expressed as:
$\frac{1}{D_{c}}=\frac{E_{c}}{L_{c}}=\sqrt{\frac{r_{c}-2m}{r_{c}^{3}}+\left(\frac{J(r_{c})+G(r_{c})\sqrt{1-\frac{2m}{r_{c}}}}{8r_{c}^{6}(r_{c}-2m)}\right)\Theta^{2}},$ (50)
@@
$J(r_{c})=152m^{3}-161m^{2}r_{c}+52mr_{c}^{2}-5r_{c}^{3},\quad G(r_{c})=29m^{2}r_{c}-27mr_{c}^{2}+3r_{c}^{3}.$ (51)
Table 1: Numerical values for the unstable $r_{\text{uns}}$ and the new stable circular orbit $r_{\text{sta}}$ of the photon sphere in NC spacetime for different $\Theta$ values, with $E=0.998$, $l=4$, and $m=1$.
@@
$r_{\text{sta}}$ (new) 2.19378 2.25867 2.31807 2.37379 2.42689
$r_{\text{uns}}$ 3.00000 2.99886 2.99739 2.99528 2.99246 2.98881
Table 1 presents numerical solutions of Eq. (32), illustrating how the stable and unstable photon sphere radii depend on $\Theta$. For $3.17\leq l\leq 4.12$ (where this range corresponds to $0\leq\Theta\leq 0.3$ and energy $E=1$), the stable circular orbit radius increases with $\Theta$, while the unstable orbit radius decreases. However, for $l>4.12$, both orbit radii increase as $\Theta$ grows.
Figure 12: Behavior of the photon sphere radius in NC spacetime. Stable and unstable circular orbits are shown as functions of $\Theta$ for fixed $l=4$, $E=1$, $m=1$ (two left panels), and as functions of $l$ for fixed $\Theta=0.3$, $E=1$, $m=1$ (two right panels).
The left panels of Fig. 12 illustrate that as $\Theta$ increases, the radius of the stable circular orbit grows, while the radius of the unstable orbit decreases. In the NC framework, the unstable photon orbit has a smaller radius than in the commutative case for $l<4.12$, whereas for $l\geq 4.12$, its radius exceeds the commutative value. This divergence arises from modifications to the gravitational field induced by non-commutativity. The right panels of Fig. 12 show the dependence of circular orbit radii on angular momentum $l$. For a fixed $\Theta=0.3$, the stable circular orbit radius decreases as $l$ increases, while the unstable orbit radius grows. This behavior highlights the interplay between angular momentum and the NC parameter in shaping the geodesic structure of spacetime.
@@ III.4.2 Lyapunov exponents
In the following, we analyze the stability and instability of the two photon spheres in NC spacetime using the Lyapunov exponent. Since proper time does not exist for null circular orbits, only the coordinate time Lyapunov exponent is considered.
We begin by computing the angular frequency and the coordinate time period, given by Eqs. (36) and (39), respectively:
@@
Next, we derive the coordinate time Lyapunov exponent for photon orbits around the NC Schwarzschild BH as:
$\hat{\lambda}_{c}^{\text{Null}}=\sqrt{-\frac{3((r_{c}-4m)(r_{c}-2m))}{r_{c}^{4}}+\left(\frac{O(r_{c})+I(r_{c})\sqrt{1-\frac{2m}{r_{c}}}}{16(r_{c}-2m)^{2}r_{c}^{7}}\right)\Theta^{2}},$ (53)
@@
Fig. 14 illustrates the behavior of the coordinate time Lyapunov exponent for photons as a function of $r_{c}$ (left panel) and $\Theta$ (right panel). The left panel shows that instability initially increases to a peak whose magnitude increases with higher $\Theta$, before decreasing as $r_{c}$ increases, eventually vanishing for $r_{c}\geq 4$. In NC spacetime, orbits near the event horizon exhibit reduced instability compared to the commutative case, whereas at larger distances, this trend reverses. Moreover, while instability in commutative spacetime begins at the event horizon, NC geometry shifts it outward, allowing for stable orbits to emerge near the event horizon. This observation aligns with the effective potential analysis for massless test particles, confirming the existence of a stable photon sphere around the NC Schwarzschild BH. Similar behavior has been reported in other NC geometries, as discussed in Ref. geostab5 , where stable photon spheres cannot form within the event horizon. The right panel examines the dependence of circular orbit radii $r_{c}$ on $\Theta$. It reveals that the orbit at $r_{c}=3$ remains unstable in NC spacetime, with instability increasing as $\Theta$ grows. In contrast, the newly emerging circular orbits near the event horizon exhibit reduced instability, which further decreases with increasing $\Theta$, eventually leading to stabilization. This suggests the existence of stable circular orbits near the event horizon in NC spacetime.
The ratio of the Lyapunov exponent to the angular frequency in coordinate time, $\frac{\hat{\lambda}_{c}^{\text{Null}}}{\hat{\Omega}_{c}^{\text{Null}}}$, defines the instability exponent of null circular orbits in NC spacetime:
$\frac{\hat{\lambda}_{c}^{\text{Null}}}{\hat{\Omega}_{c}^{\text{Null}}}=\sqrt{\frac{3(4m-r_{c})}{r_{c}}+\left(\frac{O(r_{c})+6J(r_{c})(8m^{2}-6mr_{c}+r_{c}^{2})+(I(r_{c})+6G(r_{c})(8m^{2}-6mr_{c}+r_{c}^{2}))\sqrt{1-\frac{2m}{r_{c}}}}{16(r_{c}-2m)^{3}r_{c}^{4}}\right)\Theta^{2}}.$ (55)
@@
Fig. 15 illustrates the variation of $\hat{\lambda}_{c}^{\text{Null}}/\hat{\Omega}_{c}^{\text{Null}}$ with $r_{c}$, exhibiting behavior similar to that in Fig. 14. Unlike in the commutative case, NC spacetime permits the existence of stable orbits near the event horizon. This contrasts with the findings of Ref. geostab5 , which confirm the stability of the newly emerging photon sphere.
Finally, the critical exponent $\hat{\gamma}$ for null circular orbits is defined as the ratio of the Lyapunov timescale, $\hat{T}_{\hat{\lambda}}=\frac{1}{\hat{\lambda}}$, to the orbital timescale, $\hat{T}_{\hat{\Omega}}=\frac{2\pi}{\hat{\Omega}}$ geostab11 ; geostab3 ; geostab12 :
$\hat{\gamma}^{\text{Null}}=\frac{\hat{T}_{\hat{\lambda}}^{\text{Null}}}{\hat{T}_{\hat{\Omega}}^{\text{Null}}}=\frac{1}{2\pi\sqrt{\frac{3(4m-r_{c})}{r_{c}}+\left(\frac{O(r_{c})+6J(r_{c})(8m^{2}-6mr_{c}+r_{c}^{2})+(I(r_{c})+6G(r_{c})(8m^{2}-6mr_{c}+r_{c}^{2}))\sqrt{1-\frac{2m}{r_{c}}}}{16(r_{c}-2m)^{3}r_{c}^{4}}\right)\Theta^{2}}}.$ (56)
@@
The numerical values of the critical exponent $\hat{\gamma}^{\text{Null}}$ for photons in various circular orbits $r_{c}$ and different values of the NC parameter $\Theta$ are summarized in Table 2. This table highlights the presence of both real and imaginary values of $\hat{\gamma}^{\text{Null}}$. Real positive values indicate unstable circular orbits geostab11 , whereas imaginary values correspond to stable circular orbits. For the photon sphere orbit in the Schwarzschild BH spacetime at $r_{c}=3$, instability persists even in the NC framework. According to the definition in Eq. (56), this implies that the Lyapunov timescale $\hat{T}_{\hat{\lambda}}$ is shorter than the orbital timescale $\hat{T}_{\hat{\Omega}}$, i.e., $\hat{T}_{\hat{\lambda}}<\hat{T}_{\hat{\Omega}}$ geostab12 . A similar behavior is observed for other values of $r_{c}$ in Table 2, where real values of $\hat{\gamma}^{\text{Null}}$ consistently signify instability. Furthermore, the impact of the NC parameter $\Theta$ on $\hat{\gamma}^{\text{Null}}$ is evident. As $\Theta$ increases, the values of $\hat{\gamma}^{\text{Null}}$ decrease for $r_{c}=3$, indicating increasing instability. The smallest values of $\hat{\gamma}^{\text{Null}}$ correspond to stronger Lyapunov instability geostab3 (see the right panel of Fig. 14). On the other hand, the imaginary values of $\hat{\gamma}^{\text{Null}}$, which denote stable circular orbits, align with the new minima in the effective potential observed for $l=4.2$ (values comparable to those in Table 1). This demonstrates the emergence of a new stable photon sphere in NC spacetime, contrasting with the instability of the external photon sphere in both commutative and NC scenarios.
III.4.3 Shadow radius and limit of the non-commutativity parameter
@@
In the equatorial plane ($\theta=\pi/2$), the radius of the BH shadow, $\hat{R}_{\text{shadow}}$, is directly related to the impact parameter $D_{c}=\bigg{|}\frac{L_{c}}{E_{c}}\bigg{|}$ by the relation $\hat{R}_{\text{shadow}}=D_{c}\big{|}_{r=r^{\text{uns}}_{\text{ps}}}$. This is significant because a light beam emitted from infinity with the critical impact parameter $D_{c}=D_{c}\big{|}_{r=r^{\text{uns}}_{\text{ps}}}$ will precisely reach the unstable circular orbit of the photon, while a light beam emitted with an impact parameter $D_{c}<D_{c}\big{|}_{r=r^{\text{uns}}_{\text{ps}}}$ will fall into the BH. The radius $r=r^{\text{uns}}_{\text{ps}}$ represents the location of the unstable photon sphere BHShadow1 ; Hassanabadi2 . This radius can be determined by solving the equation $\frac{dV_{\text{eff}}}{dr}=0$. To leading order in the NC parameter $\Theta$ and the BH mass $m$, the solution is given by:
$\hat{r}^{\text{uns}}_{\text{ps}}=3m+\left(\frac{15+74\sqrt{3}}{216m}\right)\Theta^{2}\,.$ (57)
@@
It is important to note that the radius of the stable photon sphere in NC spacetime can only be obtained numerically (see Table 1). Using the expression (57) together with (50), the shadow radius of the NC Schwarzschild BH can be expressed using $\hat{R}_{\text{shadow}}=D_{c}\big{|}_{r=r^{\text{uns}}_{\text{ps}}}$:
@@
$\hat{R}_{\text{shadow}}=\Bigg{|}\frac{L_{c}}{E_{c}}\Bigg{|}_{r=r^{\text{uns}}_{\text{ps}}}=3\sqrt{3}m\left(1+\frac{(25\sqrt{3}-2)}{432m^{2}}\Theta^{2}\right)\,,$ (58)
@@
Fig. 16 illustrates the shadow of the NC Schwarzschild BH for different values of the NC parameter $\Theta$ (left panel) and the BH mass $m$ (right panel). As shown in the left panel, the shadow radius increases with increasing $\Theta$. Similarly, the right panel indicates that for a fixed NC parameter, the shadow radius grows as the BH mass $m$ increases. This behavior suggests that the effects of non-commutativity and BH mass on the shadow radius are analogous. The NC parameter influences the spacetime geometry in a manner that mimics an increase in mass, effectively strengthening the gravitational field. This conclusion is consistent with our earlier results abdellah1 . Furthermore, we observe that the shadow radius of the NC Schwarzschild BH is larger and exhibits a slight increase with increasing $\Theta$, in contrast to the findings in Hassanabadi2 , where a significant increase with increasing $\Theta$ was reported.
@@
Armed with the above result and considering the constraint on the NC parameter namely, that the NC correction must be smaller than the precision of the experimental measurement we choose to implement the NC shadow radius of the Schwarzschild BH. Specifically, we consider a typical primordial BH at the final stage of inflation with a mass of $GM\sim 5\times 10^{-5}m$ karimabadi1 and a volume of $r\sim 1.5\times 10^{-4}m$. Moreover, the precision of the experimental measurement, according to ehc , is approximately $\sim 0.1$. Consequently, the upper limit of the NC parameter can be determined as:
$\Bigg{|}\frac{R_{\text{shadow}}-\hat{R}_{\text{shadow}}}{R_{\text{shadow}}}\Bigg{|}=\frac{(25\sqrt{3}-2)}{432m^{2}}\Theta^{2}\leq 0.1\,.$ (59)
@@
According to Ref. PLANCK1 , the physical distance measured at any time after inflation can be obtained by multiplying our results by the square of the cosmological scale factor, $a=10^{-29}$. This follows from our choice of the space-space NC matrix. By applying the above constraint to our system, we find that the NC parameter satisfies:
$\Theta^{\text{Phy}}=\sqrt{a^{2}\Theta^{2}}\leq 5.11\times 10^{-34}\,\text{m}.$ (60)
@@
It is evident that the obtained result is close to the Planck scale and is consistent with the bounds derived in our previous works, where similar constraints were obtained through experimental tests of gravity in the context of the NC gauge theory of gravity abdellah1 ; abdellah4 or by employing Bopp’s shift transformation larbi1 , and even for the NC geometry in flat spacetime mirza ; rome1 . These studies suggest a bound in the range of $(10^{-34}-10^{-31})$ m.
Moreover, we can obtain another constraint on the NC parameter using the unstable photon sphere, based on the same system described above. In this case, the lower bound of the NC parameter, derived from the unstable photon sphere with the accuracy of the Event Horizon Telescope ehc , is given by:
$\Bigg{|}\frac{r^{\text{uns}}_{\text{ps}}-\hat{r}^{\text{uns}}_{\text{ps}}}{r^{\text{uns}}_{\text{ps}}}\Bigg{|}=\left(\frac{15+74\sqrt{3}}{648m^{2}}\right)\Theta^{2}\leq 0.1\,.$ (61)
and that lead to the new constrain:
$\Theta^{\text{Phy}}\leq 3.36\times 10^{-34}\,\text{m}.$ (62)
As we can see, this constraint is of the same order as the one obtained from the shadow radius of this BH. The difference arises due to the unstable photon sphere being closer to the NC event horizon compared to the shadow radius.
Furthermore, it is possible to generalize this application to examine the constraint on the NC parameter from the innermost stable circular orbit (ISCO) of a massive test particle and to compare it with the constraint obtained from the photon orbit. The outer ISCO can be determined using the effective potential with $h=1$ and $E=c=1$, together with $l_{crit}=3.5m$. By solving the following condition, $\frac{dV_{\text{eff}}}{dr}=0$, we obtain:
$\hat{r}^{\text{ISCO}}_{\text{outer}}\sim 7m+\left(\frac{0.0677306}{m}\right)\Theta^{2}\sim 7m+\left(\frac{\sqrt{17}}{61m}\right)\Theta^{2}\,.$ (63)
We assume that the observation of any close orbit of a massive or massless test particle is subject to the same accuracy constraint, which is approximately $10\%$ ehc . Consequently, the constraint on $\Theta$ from the ISCO of a massive particle is given by:
$\Bigg{|}\frac{r^{\text{ISCO}}-\hat{r}^{\text{ISCO}}_{\text{outer}}}{r^{\text{ISCO}}}\Bigg{|}=\left(\frac{\sqrt{17}}{427m^{2}}\right)\Theta^{2}\leq 0.1\,.$ (64)
which give us:
$\Theta^{\text{Phy}}\leq 1.61\times 10^{-33}\,\text{m}.$ (65)
It is clear that the constraint obtained from the trajectory of a massive particle is larger than that obtained from the photon trajectory, which aligns with our previous findings in Ref. abdellah4 . Notably, the inner ISCO and the stable photon sphere are unique predictions of NC spacetime and can only be determined numerically. Any constraint derived from these new orbits requires direct observational data to obtain a numerical resolution that allows for a meaningful comparison with experimental observations. We believe that using these two new circular orbits provides a better estimation of the NC parameter compared to the outer orbits.
It is particularly noteworthy that as we move closer to the event horizon of the BH, the NC parameter is constrained to smaller values. As demonstrated in the previous comparison between the unstable photon sphere and the shadow radius, where $\hat{r}^{\text{uns}}_{\text{ps}}<\hat{R}_{\text{shadow}}$ leads to a smaller value of $\Theta$. Thus, since the new circular orbits are closer to the event horizon, the constraint on the NC parameter derived from the inner ISCO or the stable photon sphere is expected to be smaller than those obtained in previous discussions, bringing it even closer to the Planck scale.
Remarkably, in this work, the estimation of the NC parameter is directly related to the accuracy of the experimental measurement of the BH shadow radius. This implies that as the precision of the measurement improves, the estimation of the NC parameter becomes more refined. A similar observation was made in Ref. abdellah4 , where variations in experimental accuracy led to different estimations of the NC parameter.